# Sensing for Free: Learn to Localize More Sources than Antennas without Pilots

### v1

@@ II-B Problem Statement
Based on $T$ snapshots of received signals at the $M$-element SLA, i.e., $\mathbf{Y}_{\bm{\Omega}}$, the target is to estimate the DOAs of $K$ sources, i.e., $\bm{\theta}$, in which $K$ could be larger than $M$. We note that the sources $\mathbf{s}(t)$ are unknown uplink data symbols, in contrast to pilot symbols that are perfectly known [41]. The multi-source localization is carried out at the same time during data transmission and thus requires no additional time or frequency resources. As the true and estimated DOAs are two sets, to remove the permutation ambiguity when comparing their distance, we define the mean square error (MSE) of the estimated angles $\hat{\bm{\theta}}$ and the ground-truth $\bm{\theta}$ as
@@
The value of the problem is that, once this is achieved, we can localize more sources than the number of physical antennas by using unknown data symbols, at the same time of data transmission, without using any additional pilots or radio resources. In this sense, sensing is achieved “for free”.
@@ III Prior Art
While there are no previous works for ISAC with sparse arrays using unknown data symbols, we summarize the state-of-the-art DL-based algorithms for sparse array DOA estimation using Gaussian signals, which will serve as important benchmarks later. The dataset is defined as $\{\mathbf{Y}_{\bm{\Omega}}^{(i)},K^{(i)},\bm{\theta}^{(i)}\}_{i=1}^{I}$, where $I$ is the number of samples in the dataset and the superscript $(\cdot)^{(i)}$ denotes the $i$-th sample. For the $i$-th data sample, the DOA estimation algorithm takes both the received signal snapshots at the SLA, i.e., $\mathbf{Y}_{\bm{\Omega}}^{(i)}\in\mathbb{C}^{M\times T}$, and the number of sources, i.e., $K^{(i)}$ as inputs, while the outputs are the estimated DOAs whose ground-truth labels are $\bm{\theta}^{(i)}\in\mathbb{R}^{K\times 1}$.
@@
DCR-G-Fro [29] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{H}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\|\hat{\mathbf{R}}-\mathbf{R}\|_{F}$
DCR-G-Aff [29] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{H}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\bigl{\|}\ln\bigl{(}\mathbf{R}^{-\tfrac{1}{2}}\,\hat{\mathbf{R}}\,\mathbf{R}^{-\tfrac{1}{2}}\bigr{)}\bigr{\|}_{F}$
DCR-T [30] $\hat{\mathbf{v}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}}),\hat{\mathbf{R}}=\operatorname{Toep}(\hat{\mathbf{v}})$ $\bigl{\|}\hat{\mathbf{v}}-[\mathbf{A}(\bm{\theta})\mathbf{A}^{H}(\bm{\theta})]_{1,:}^{T}\bigr{\|}_{2}$
SRL [31] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{H}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\mathrm{dist}_{\mathrm{sub}}\bigl{(}\hat{\mathbf{R}},\mathbf{R}\bigr{)}$
@@ III-A Covariance Reconstruction
Building upon the intuition that the noiseless SCM of the ULA should be both Hermitian and positive semi-definite, the authors of [29] trained a CNN to estimate an auxiliary matrix $f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})\in\mathbb{C}^{N\times N}$, whose Gramian $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{H}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ is an estimate of $\mathbf{R}$, which will be used in subspace methods. To measure the performance of deep covariance reconstruction (DCR), the authors have proposed two loss functions, i.e., the Frobenius norm and the affine invariant distance, called DCR-G-Fro and DCR-G-Aff, respectively.
Noticing that the SCM of ULA is a Toeplitz matrix characterized by its first row, the authors of [30] proposed to estimate it using a CNN $f_{\mathcal{W}}(\cdot)$ whose output is $\hat{\mathbf{\mathbf{v}}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})\in\mathbb{C}^{N\times 1}$. The loss function is chosen as the Euclidean norm between $\hat{\mathbf{\mathbf{v}}}$ and $[\mathbf{A}(\bm{\theta})\mathbf{A}^{H}(\bm{\theta})]_{1,:}^{T}$. The estimated SCM is then constructed as a Toeplitz matrix according to the estimated first row, i.e., $\hat{\mathbf{R}}=\operatorname{Toep}(\hat{\mathbf{v}})$, which will be used in subspace methods. This algorithm is shortened as DCR-T.
The authors of [31] observed that subspace methods require only that the reconstructed SCM $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{H}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ can share the same signal subspace as the noiseless ULA SCM $\mathbf{R}$. Exact recovery of $\mathbf{R}$ is a harder problem that is not necessary. Based on this, they defined various metrics that instead measure the distances between the signal subspaces of $\hat{\mathbf{R}}$ and $\mathbf{R}$, denoted by $\operatorname{dist_{sub}}(\hat{\mathbf{R}},\mathbf{R})$. We denote the eigen-decomposition of $\hat{\mathbf{R}}$ as
$\hat{\mathbf{R}}=\begin{bmatrix}\hat{\mathbf{U}}&\hat{\mathbf{V}}\end{bmatrix}\begin{bmatrix}\hat{\bm{\Lambda}}_{K}\\ &\hat{\bm{\Lambda}}_{N-K}\end{bmatrix}\begin{bmatrix}\hat{\mathbf{U}}^{H}\\[3.00003pt] \hat{\mathbf{V}}^{H}\end{bmatrix},$ (4)
where $\hat{\bm{\Lambda}}_{K}$ and $\hat{\bm{\Lambda}}_{N-K}$ are both diagonal matrices consisting of the $K$ largest and the $(N-K)$ smallest eigen-values. The columns of $\hat{\mathbf{U}}\in\mathbb{C}^{N\times K}$ and $\hat{\mathbf{V}}\in\mathbb{C}^{N\times(N-K)}$ consist of the eigen-vectors that correspond to the signal and the noise subspaces, respectively. A similar decomposition can be performed for the noiseless ULA SCM $\mathbf{R}$, which will yield $\mathbf{U}$ and $\mathbf{V}$. The signal subspaces for the ground-truth and the estimated ULA SCMs are the range spaces of $\mathbf{U}$ and $\hat{\mathbf{U}}$, respectively. The authors used principle angles as a measure of the distance between signal subspaces [32], given by
$\operatorname{dist_{sub}(\hat{\mathbf{R}},\mathbf{R})}=\bigl{\lVert}\cos^{-1}\bigl{(}\sigma\bigl{(}\mathbf{U}^{H}\hat{\mathbf{U}}\bigr{)}\bigr{)}\bigr{\rVert}_{2},$ (5)
@@ III-B Subspace Methods
$\mathbf{a}^{H}(\theta_{k})\hat{\mathbf{V}}\hat{\mathbf{V}}^{H}\mathbf{a}(\theta_{k})=\lVert\hat{\mathbf{V}}^{H}\mathbf{a}(\theta_{k})\rVert_{2}^{2}=0,$ (6)
@@
III-B1 MUSIC
For each candidate angle $\theta_{k}$, the corresponding array response vector $\mathbf{a}({\theta}_{k})$ is projected onto the noise subspace, and the resulting MUSIC spectrum is defined as
$P_{\text{MUSIC}}(\theta_{k})=\frac{1}{\lVert\hat{\mathbf{V}}^{H}\mathbf{a}(\theta_{k})\rVert_{2}^{2}},$ (7)
which is evaluated on a dense grid of $G$ points. A peak search of the spectrum identifies $K$ estimated angles $\hat{\bm{\theta}}$. The dominant computational cost is the eigen-decomposition with $O(N^{3})$. In addition, evaluating the MUSIC spectrum on the $G$ grid points costs a complexity of $O(N^{2}G)$, in which $G$ can be quite large. For an SLA with $M$ antennas, its corresponding virtual ULA contains $N$ elements much larger than $M$ [14], i.e., $N\gg M$. For example, for a minimum redundancy SLA [13] with $M=28$ antennas, the corresponding $N=244$ [43].
III-B2 Root-MUSIC
As an alternative, Root-MUSIC algorithm seeks to find the roots of (6), which are related to DOAs. The left hand side of (6) is a conjugate reciprocal polynomial of $z=e^{j\pi\sin(\theta)}$, given by
$P(z)=\sum_{m=-(N-1)}^{N-1}c_{m}z^{-m},$ (8)
in which $c_{m}=c_{-m}^{*}=\sum_{i-j=m}[\hat{\mathbf{V}}\hat{\mathbf{V}}^{H}]_{i,j}$, with $m=-(N-1),\ldots,N-1$ [24]. The polynomial $P(z)$ has $N-1$ reciprocal pairs of zeros. The $K$ roots lying inside and being the closest to the unit circle are related to the estimated DOAs. The $k$-th estimated DOA is obtained by $\hat{\theta}_{k}=\sin^{-1}(\frac{\arg(z_{k})}{\pi})$. Similar to MUSIC, the dominant complexity is $O(N^{3})$ of the eigen-decomposition step. In addition, the root finding process using companion matrix also costs $O(N^{3})$ complexity.
@@ IV Our Proposed Snap-TF Algorithm
4:  Initialize: Initial feature set $\mathbf{S}^{(0)}\leftarrow[\Re(\mathbf{Y}_{\bm{\Omega}}^{T}),\Im(\mathbf{Y}_{\bm{\Omega}}^{T})]^{T}\in\!\mathbb{R}^{2M\times T}$
@@
6:     Self-Attention Block:
7:      $\mathbf{Z}^{(\ell)}\leftarrow\mathrm{LayerNorm}\mkern 1.0mu\Bigl{(}\mathbf{S}^{(\ell-1)}+{\smash{\overbrace{(\mathbf{W}_{V}^{(\ell)}\mathbf{S}^{(\ell-1)})}^{\scriptstyle\mathbf{V}^{(\ell)}\in\,\mathbb{R}^{d_{\text{attn}}\times T}}}\cdot\mathrm{Softmax}\mkern 1.0mu\Bigl{(}\frac{\smash{\overbrace{(\mathbf{W}_{K}^{(\ell)}\mathbf{S}^{(\ell-1)})^{T}}^{\scriptstyle(\mathbf{K}^{(\ell)})^{T}\in\,\mathbb{R}^{T\times d_{\text{attn}}}}}\cdot\smash{\overbrace{(\mathbf{W}_{Q}^{(\ell)}\mathbf{S}^{(\ell-1)})}^{\scriptstyle\mathbf{Q}^{(\ell)}\in\,\mathbb{R}^{d_{\text{attn}}\times T}}}}{\sqrt{d_{\text{attn}}}}}\Bigr{)}$
8:     Feed-Forward Block:
9:      $\mathbf{S}^{(\ell)}\leftarrow\mathrm{LayerNorm}\mkern 1.0mu\bigl{(}\mathbf{Z}^{(\ell)}+\mathbf{W}_{2}^{(\ell)}\,\mathrm{ReLU}\mkern 1.0mu\bigl{(}\mathbf{W}_{1}^{(\ell)}\mathbf{Z}^{(\ell)}+\mathbf{b}_{1}^{(\ell)}\mathbf{1}^{T}\bigr{)}+\mathbf{b}_{2}^{(\ell)}\mathbf{1}^{T}\bigr{)}$
@@ IV-A1 Permutation Invariance
The algorithm should be permutation invariant to the order of the snapshots. This is because the received signals $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ are drawn i.i.d. across different snapshot $t$, as the DOAs are static during estimation and the symbols and noise vectors are both i.i.d. across different snapshots. This means that the order of the snapshots should not affect the estimated DOAs. Denote the proposed Snap-TF DOA estimation network as $g_{\mathbb{\mathcal{W}}}(\mathbf{\mathbf{Y}_{\bm{\Omega}}})$, where $\mathcal{W}$ is the set of trainable parameters. By permutation invariance, we refer to the property that the permutating the columns of $\mathbf{Y}_{\bm{\Omega}}$, i.e., the order of snapshots, will not change the output DOA estimates of the neural network, i.e., $g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}})=g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}}\bm{\Pi}_{T}),\forall\bm{\Pi}_{T}\in\mathcal{P}_{T}$, where $\mathcal{P}_{T}$ here denotes the set of all $T\times T$ permutation matrices. Otherwise, the network will learn misleading information regarding the order of the snapshots, and have poor generalization performance.
Existing DL-augmented subspace methods introduced in the last section inherently satisfy this requirement since their input is the sample SCM $\tilde{\mathbf{R}}_{\bm{\Omega}}\in\mathbb{C}^{M\times M}$ whose dimension remains constant regardless of snapshot count $T$, and whose value is invariant with respect to the order of snapshots $\mathbf{y}_{\bm{\Omega}}(t)$. However, designing end-to-end neural networks that can process raw signal snapshots to capture higher-order statistics while maintaining similar generalization capabilities is much more challenging, and requires careful architectural considerations. Popular feed-forward and recurrent neural networks (RNNs) are sensitive to the input order and hence not suitable.
IV-A2 Generalization Across Varying Snapshot Counts
In real deployments the snapshot count $T$ is rarely fixed. It shrinks when coherence times are short or processing budgets are tight, and it grows when the coherence times are longer. A robust DOA estimator should therefore generalize across arbitrary $T$, returning accurate estimates whether it sees fewer or more snapshots than in training. The natural way to meet this requirement is to treat the received signal snapshots $\mathbf{Y}_{\bm{\Omega}}\in\mathbb{C}^{M\times T}$ as an unordered set, with each column $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ being an element of the set whose order is exchangeable with other set elements. Consequently, the network must be permutation invariant and able to process the input of different snapshot counts without retraining.

### v2

@@ II-B Problem Statement
Based on $T$ snapshots of received signals at the $M$-element SLA, i.e., $\mathbf{Y}_{\bm{\Omega}}$, the target is to estimate the DOAs of $K$ sources, i.e., $\bm{\theta}$, in which $K$ could be larger than $M$. We work under the “sensing for free” setup (see Remark 1). As the true and estimated DOAs are two sets, to remove the permutation ambiguity when comparing their distance, we define the mean square error (MSE) of the estimated angles $\hat{\bm{\theta}}$ and the ground-truth $\bm{\theta}$ as
@@ III Prior Art
While there are no previous works for multi-source localization with sparse arrays using unknown data symbols, we summarize the state-of-the-art DL-based algorithms for sparse array DOA estimation using Gaussian signals, which will serve as important benchmarks later. The dataset is defined as $\{\mathbf{Y}_{\bm{\Omega}}^{(i)},K^{(i)},\bm{\theta}^{(i)}\}_{i=1}^{I}$, where $I$ is the number of samples in the dataset and the superscript $(\cdot)^{(i)}$ denotes the $i$-th sample. For the $i$-th data sample, the DOA estimation algorithm takes both the received signal snapshots at the SLA, i.e., $\mathbf{Y}_{\bm{\Omega}}^{(i)}\in\mathbb{C}^{M\times T}$, and the number of sources, i.e., $K^{(i)}$ as inputs, while the outputs are the estimated DOAs whose ground-truth labels are $\bm{\theta}^{(i)}\in\mathbb{R}^{K\times 1}$.
@@
DCR-G-Fro [32] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{\mathsf{H}}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\|\hat{\mathbf{R}}-\mathbf{R}\|_{F}$
DCR-G-Aff [32] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{\mathsf{H}}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\bigl\|\ln\bigl(\mathbf{R}^{-\tfrac{1}{2}}\,\hat{\mathbf{R}}\,\mathbf{R}^{-\tfrac{1}{2}}\bigr)\bigr\|_{F}$
[…]
SRL [34] $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{\mathsf{H}}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ $\mathrm{dist}_{\mathrm{sub}}\bigl(\hat{\mathbf{R}},\mathbf{R}\bigr)$
@@ III-A Covariance Reconstruction
Building upon the intuition that the noiseless SCM of the ULA should be both Hermitian and positive semi-definite, the authors of [32] trained a CNN to estimate an auxiliary matrix $f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})\in\mathbb{C}^{N\times N}$, whose Gramian $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{\mathsf{H}}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ is an estimate of $\mathbf{R}$, which will be used in subspace methods. To measure the performance of deep covariance reconstruction (DCR), the authors have proposed two loss functions, i.e., the Frobenius norm and the affine invariant distance, called DCR-G-Fro and DCR-G-Aff, respectively.
[…]
The authors of [34] observed that subspace methods require only that the reconstructed SCM $\hat{\mathbf{R}}=f_{\mathcal{W}}(\tilde{\mathbf{R}}_{\bm{\Omega}})f_{\mathcal{W}}^{\mathsf{H}}(\tilde{\mathbf{R}}_{\bm{\Omega}})$ can share the same signal subspace as the noiseless ULA SCM $\mathbf{R}$. Exact recovery of $\mathbf{R}$ is a harder problem that is not necessary. Based on this, they defined various metrics that instead measure the distances between the signal subspaces of $\hat{\mathbf{R}}$ and $\mathbf{R}$, denoted by $\operatorname{dist_{sub}}(\hat{\mathbf{R}},\mathbf{R})$. We denote the eigen-decomposition of $\hat{\mathbf{R}}$ as
$\hat{\mathbf{R}}=\begin{bmatrix}\hat{\mathbf{U}}&\hat{\mathbf{V}}\end{bmatrix}\begin{bmatrix}\hat{\bm{\Lambda}}_{K}\\ &\hat{\bm{\Lambda}}_{N-K}\end{bmatrix}\begin{bmatrix}\hat{\mathbf{U}}^{\mathsf{H}}\\[3.00003pt] \hat{\mathbf{V}}^{\mathsf{H}}\end{bmatrix}$, (4)
where $\hat{\bm{\Lambda}}_{K}$ and $\hat{\bm{\Lambda}}_{N-K}$ are both diagonal matrices consisting of the $K$ largest and the (NK)(N-K) smallest eigen-values. The columns of $\hat{\mathbf{U}}\in\mathbb{C}^{N\times K}$ and $\hat{\mathbf{V}}\in\mathbb{C}^{N\times(N-K)}$ consist of the eigen-vectors that correspond to the signal and the noise subspaces, respectively. A similar decomposition can be performed for the noiseless ULA SCM $\mathbf{R}$, which will yield $\mathbf{U}$ and $\mathbf{V}$. The signal subspaces for the ground-truth and the estimated ULA SCMs are the range spaces of $\mathbf{U}$ and $\hat{\mathbf{U}}$, respectively. The authors used principle angles as a measure of the distance between signal subspaces [35], given by
$\operatorname{dist_{sub}(\hat{\mathbf{R}},\mathbf{R})}=\bigl\lVert\cos^{-1}\bigl(\sigma\bigl(\mathbf{U}^{\mathsf{H}}\hat{\mathbf{U}}\bigr)\bigr)\bigr\rVert_{2},$ (5)
@@ III-B Subspace Methods
$\mathbf{a}^{\mathsf{H}}(\theta_{k})\hat{\mathbf{V}}\hat{\mathbf{V}}^{\mathsf{H}}\mathbf{a}(\theta_{k})=\lVert\hat{\mathbf{V}}^{\mathsf{H}}\mathbf{a}(\theta_{k})\rVert_{2}^{2}=0,$ (6)
@@ IV Our Proposed Snap-TF Algorithm
4:Initialize: Initial feature set $\mathbf{S}^{(0)}\leftarrow[\Re(\mathbf{Y}_{\bm{\Omega}}^{\mathsf{T}}),\Im(\mathbf{Y}_{\bm{\Omega}}^{\mathsf{T}})]^{\mathsf{T}}\in\!\mathbb{R}^{2M\times T}$
@@
6:  Self‑Attention Block:
7:   $\mathbf{Z}^{(\ell)}\leftarrow\mathrm{LayerNorm}\mkern 1.0mu\Bigl(\mathbf{S}^{(\ell-1)}+{\smash{\overbrace{(\mathbf{W}_{V}^{(\ell)}\mathbf{S}^{(\ell-1)})}^{\scriptstyle\mathbf{V}^{(\ell)}\in\,\mathbb{R}^{d_{\text{attn}}\times T}}}\cdot\mathrm{Softmax}\mkern 1.0mu\Bigl(\frac{\smash{\overbrace{(\mathbf{W}_{K}^{(\ell)}\mathbf{S}^{(\ell-1)})^{\mathsf{T}}}^{\scriptstyle(\mathbf{K}^{(\ell)})^{\mathsf{T}}\in\,\mathbb{R}^{T\times d_{\text{attn}}}}}\cdot\smash{\overbrace{(\mathbf{W}_{Q}^{(\ell)}\mathbf{S}^{(\ell-1)})}^{\scriptstyle\mathbf{Q}^{(\ell)}\in\,\mathbb{R}^{d_{\text{attn}}\times T}}}}{\sqrt{d_{\text{attn}}}}}\Bigr)$
8:  Feed‑Forward Block:
9:   $\mathbf{S}^{(\ell)}\leftarrow\mathrm{LayerNorm}\mkern 1.0mu\bigl(\mathbf{Z}^{(\ell)}+\mathbf{W}_{2}^{(\ell)}\,\mathrm{ReLU}\mkern 1.0mu\bigl(\mathbf{W}_{1}^{(\ell)}\mathbf{Z}^{(\ell)}+\mathbf{b}_{1}^{(\ell)}\mathbf{1}^{\mathsf{T}}\bigr)+\mathbf{b}_{2}^{(\ell)}\mathbf{1}^{\mathsf{T}}\bigr)$
@@ IV-A1 Permutation Invariance
The algorithm should be permutation invariant to the order of the snapshots. This is because the received signals $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ are drawn i.i.d. across different snapshot $t$, as the DOAs are static during estimation and the symbols and noise vectors are both i.i.d. across different snapshots. This means that the order of the snapshots should not affect the estimated DOAs. Denote the proposed Snap-TF DOA estimation network as $g_{\mathbb{\mathcal{W}}}(\mathbf{\mathbf{Y}_{\bm{\Omega}}})$, where $\mathcal{W}$ is the set of trainable parameters. By permutation invariance, we refer to the property that permuting the columns of $\mathbf{Y}_{\bm{\Omega}}$ (i.e., a permutation of the snapshot index) will not change the output DOA estimates of the neural network, i.e., $g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}})=g_{\mathbb{\mathcal{W}}}(\mathbf{Y}_{\bm{\Omega}}\bm{\Pi}_{T}),\forall\bm{\Pi}_{T}\in\mathcal{P}_{T}$, where $\mathcal{P}_{T}$ denotes the set of all $T\times T$ permutation matrices. Otherwise, the network will learn misleading information regarding the order of the snapshots, and have poor generalization performance.
Existing DL-augmented subspace methods introduced in the last section inherently satisfy this requirement since their input is the sample SCM $\tilde{\mathbf{R}}_{\bm{\Omega}}\in\mathbb{C}^{M\times M}$ whose dimension remains constant regardless of the number of snapshots $T$, and whose value is invariant with respect to a permutation of the snapshot index $t$ in $\mathbf{y}_{\bm{\Omega}}(t)$. However, designing end-to-end neural networks that can process raw signal snapshots to capture higher-order statistics while maintaining similar generalization capabilities is much more challenging, and requires careful architectural considerations. Popular feed-forward and recurrent neural networks (RNNs) are sensitive to the input order and hence not suitable for the considered problem.
IV-A2 Generalization to Varying Numbers of Snapshots
In real-world deployments the number of snapshots $T$ is rarely fixed. It shrinks when coherence times are short or processing budgets are tight, and it grows when the coherence times are longer. A robust DOA estimator should therefore generalize across arbitrary $T$, returning accurate estimates whether it sees fewer or more snapshots than in training. The natural way to meet this requirement is to treat the received signal snapshots $\mathbf{Y}_{\bm{\Omega}}\in\mathbb{C}^{M\times T}$ as an unordered set, with each column $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$ being an element of the set whose order is exchangeable with other set elements. Consequently, the network must be permutation invariant and able to process the input of different numbers of snapshots without retraining.